# From Bugs to Benchmarks: A Comprehensive Survey of Software Defect Datasets

### v1

@@ 6. Uses of Software Defect Datasets
Understanding how software defect datasets are used in practice is essential not only for assessing their impact on software engineering research and practice, but also for revealing current needs in software engineering and beyond. For a larger picture on the uses of software defect datasets, we have conducted a systematic analysis of research work citing the top five most cited [^4] As of January 1, 2025, on Google Scholar. software defect dataset papers among the ones considered in this survey, which are Defects4J (Just et al., 2014), Bug Prediction Dataset (D’Ambros et al., 2012), Big-Vul (Fan et al., 2020), the ManyBugs and IntroClass benchmarks (Le Goues et al., 2015), and SWE-Bench (Jimenez et al., 2024). All five datasets are general-purpose and not limited to any specific application domain. Four datasets primarily focus on functional defects, while Big-Vul specifically targets security vulnerabilities. In terms of programming language coverage, Defects4J and the Bug Prediction Dataset collect defects from Java projects; Big-Vul and the ManyBugs and IntroClass benchmarks focus on C/C++ defects; and SWE-Bench, the most recent among them, collects defects from Python projects.
@@
We categorize the uses of software defect datasets by conducting a systematic analysis of papers citing these datasets. Initially, we manually review a subset of citing papers to develop a preliminary categorization. Using an open-coding approach, we continuously refine the categories as we review additional papers, leading to the following categorization scheme:
@@
It should be noted that the categories of empirical research and technical evaluations are not mutually exclusive, as a paper may employ software defect datasets for both purposes. For instance, a paper that empirically evaluates existing program repair tools using established software defect datasets is classified into both empirical research and the repair category of technical evaluations. In contrast, a paper that introduces a new repair technique and evaluates it using defect datasets falls solely under technical evaluation. Based on this scheme, we employed LLMs to automate the categorization of the remaining papers. Lastly, we manually validated the LLMs’ classifications and found them to be accurate and consistent.
@@
As illustrated in Figure 9(a), empirical research and technical evaluations are the two dominant categories, together accounting for the majority of software defect dataset uses. Technical evaluations, which can be further divided into evaluations of detection, testing, and repair techniques, demonstrate a growing demand for software defect datasets, as shown in Figure 9(b). This section focuses on these two primary categories to provide a comprehensive overview of the uses of software defect datasets. As our analysis specifically focuses on direct dataset utilization, we exclude papers that introduce new benchmarks and the “others” category from the detailed discussion.
6.1. Empirical Research
Software defect datasets are widely utilized in empirical research to understand various aspects of software defects. Such empirical research can be broadly categorized into two groups. On one hand, studies may focus directly on analyzing software defects themselves, typically with regard to their manifestations, symptoms, root causes, and fixes. For example, Sobreira et al. (2018) present an in-depth analysis of 395 patches from Defects4J. By analyzing the patch size, location, and the underlying repair patterns, the study provides insights and future direction for automated program repair, such as highlighting the necessity for repair strategies capable of effectively handling method calls. On the other hand, empirical research can also explore interactions between software defects and existing software engineering techniques. For example, Durieux et al. (2019) conduct an empirical evaluation on eleven automated program repair tools using five software defect datasets. Their study shows that the effectiveness of automated program repair tools varies significantly across different datasets and concludes various causes of failures for the automated program repair. Both types of empirical research significantly benefit from pre-collected, well-structured software defect datasets, as these datasets offer diverse, realistic, and accessible instances of software defects, thus enabling rigorous analysis and reliable comparisons across different studies.
@@ 6.2. Technical Evaluations
Another primary use of software defect datasets is to evaluate existing or new software engineering techniques, such as fault detection/localization (e.g., (Sarhan et al., 2022; An and Yoo, 2022; Schlichtig et al., 2023; Zeng et al., 2022; Yan et al., 2022; Wen et al., 2021; Wang et al., 2021b; Yan et al., 2021; Sarhan et al., 2021; Perez et al., 2021; Kim et al., 2021a; Nilizadeh et al., 2024; Sun et al., 2024)), test generation/prioritization (e.g., (Zhang et al., 2024e; Jabbar et al., 2023; Rotaru and Vescan, 2023; Chaim et al., 2023; Chen et al., 2023a; Almulla and Gay, 2022; Biswas et al., 2022; Gazzola et al., 2022)), and automated program repair (e.g., (Ye et al., 2023b; Liu et al., 2023; Tian et al., 2023a; Fan et al., 2023; Chakraborty et al., 2022)). Fault detection/localization aims to identify defects in software in an effective, efficient, and automated manner. Test generation/prioritization is about generating or prioritizing test cases to maximize the likelihood of detecting defects, while minimizing the cost of testing. Both detection and testing techniques have long attracted significant interest from the software engineering, programming languages, and other communities, as such techniques are essential for improving software quality and reliability. The goal of automated program repair (Le Goues et al., 2019) is to fix defective software without human intervention, which is a challenging task that has gained increasing attention in recent years. The evaluation of these techniques is crucial for understanding their effectiveness, efficiency, and limitations, and software defect datasets play a critical role in this process. By providing standardized, realistic, and reproducible instances of software defects, software defect datasets enable researchers and practitioners to evaluate their techniques in a controlled environment.
@@
Figure 9(a) illustrates that detection and repair techniques are the most frequently evaluated. Defects4J is used for evaluating all three types of techniques, and it is also the predominant dataset used for evaluating testing techniques. The Bug Prediction Dataset (D’Ambros et al., 2012) and Big-Vul (Fan et al., 2020) are primarily used for evaluating detection techniques, while both the ManyBugs and IntroClass benchmarks (Le Goues et al., 2015) and SWE-Bench (Jimenez et al., 2024) are used more extensively for evaluating repair techniques. Figure 9(b) illustrates the unprecedented growth in the number of proposed software engineering techniques, particularly in the area of automated program repair. This surge in recent years underscores the increasing importance and demand for high-quality software defect datasets.
@@ 7.1. Creation of Software Defect Datasets
Though various methodologies for constructing defect datasets have been proposed (Section 4), the balance between scale, diversity, quality, and cost remains a challenge. Intuitively, datasets of high quality tend to be smaller and more costly to create, such as the hand-crafted Defects4J (Just et al., 2014). In this sense, further efforts are still needed for fully automated defect mining. However, an important bottleneck for automated mining is the lack of a ground truth that indicates the actual defects present in the codebase. For example, keyword-based mining, the most commonly used technique, shows a relatively low precision of 51.9% (Tian et al., 2012). It would be desirable to establish a ground truth without manual labelling or testing. BugSwarm (Tomassi et al., 2019) and Bears (Madeiral et al., 2019) have shown that by leveraging the fail-pass pattern in CI/CD pipelines, one can automatically mine defects with high precision. Such techniques may be adopted into the mining processes for other sources, without narrowing to CI/CD build histories. Similarly, ReposVul (Wang et al., 2024) have shown that the combination of LLMs and static analysis can achieve up to 95% precision in defect mining. Such approaches can be further explored to improve the quality of defect datasets while keeping the manual effort low.
@@
Though tremendous efforts have been made in defect mining, the granularity of defect datasets is usually at the repository-, file-, or method-level, without fully precise defect location information that pinpoints defects, e.g., at the line-level. Ideally, datasets should isolate defects, e.g., by removing any unrelated code changes from a bug-fixing commit. For example, to isolate actual defects from newly added features, refactorings, and other code changes, Defects4J (Just et al., 2014) has manually reviewed and filtered the code diffs that fix a given bug. Such a manual review and filtering process is costly and time-consuming, though; thus, it is preferable to have automated techniques to precisely locate defects. Such isolation techniques could be based on refactoring detection (Jiang et al., 2021) or probabilistic delta debugging (Song et al., 2023). Another related problem is the annotation of defects, e.g., with the defect’s category. Though some datasets provide such annotations, they are typically manually labelled or inferred from other metadata such as exceptions (Tomassi et al., 2019). Recently, machine learning and natural language processing techniques have shown promising results for automated defect annotation (Song et al., 2023; Xu et al., 2019; Ahmed et al., 2024). More investigation is needed on the effectiveness and efficiency of such techniques, and on how they can be integrated into the process of creating defect datasets.
@@
Effectively organizing a dataset is crucial for maximizing its availability and usability. Currently, there is no standardized format for organizing software defect datasets, which can lead to difficulties in dataset access and usage. GitHub is the most popular platform for hosting datasets and tool chains, but it does not ensure long-term dataset availability and is susceptible to data loss or deletion. It is desirable to have datasets archived in a more stable and reliable repository, such as Zenodo or figshare, in addition to GitHub (Association for Computing Machinery (2025), ACM). Furthermore, the organization of datasets should facilitate future evaluation of new techniques. Section 5.2 shows that half of the datasets provide only textual metadata (i.e., meta-level), which is insufficient for running software defects. Future efforts should focus on providing more comprehensive software artifacts for defects, such as source code, build scripts, test cases, or even the original execution environment.
@@ 7.2. Maintenance of Software Defect Datasets
Ensuring the reproducibility of defects in software defect datasets is crucial for researchers to validate their findings and advance existing work. However, achieving reproducibility poses a significant challenge in software defect datasets, regardless of the construction methodology or dataset organization (Zhu and Rubio-González, 2023). Recent work shows that the major reason for defects not being reproducible is inaccessible dependencies, which datasets can mitigate by caching dependencies (Zhu and Rubio-González, 2023; Saavedra et al., 2024; Silva et al., 2024). In this direction, more techniques are required to improve the long-term reproducibility of datasets while maintaining the balance between dataset size and usability.
@@
Software defect datasets can greatly benefit from further evolution after their initial publication. However, currently the majority of datasets are not updated or expanded after their initial release, leading to outdated information. Specifically, datasets that do not evolve may lead to data leakage when evaluating machine learning-based approaches, especially for LLMs, as the defects may be in the training corpus (Tian et al., 2023b). To address this issue, future datasets could support continuous updates by adding new defects while maintaining historical data for longitudinal studies. Continuous mining with automated pipelines (Tomassi et al., 2019; Zhu et al., 2023) has shown its potential to keep datasets up-to-date, and community-driven efforts represent a promising direction for dataset evolution (McGuire et al., 2023). Recent studies have demonstrated that LLMs can be utilized to generate new defects (Xia et al., 2024). Further research is needed to explore best practices for dataset evolution, including how to balance the addition of new defects with the maintenance of historical data, how to ensure the quality and relevance of updated datasets, and how to incentivize dataset creators to keep evolving a published dataset.
@@ 8. Conclusion
This survey provides a comprehensive overview of the landscape of software defect datasets, including their scope, construction, availability, usability, and practical uses. As fundamental resources to the software engineering, programming languages, computer security, computer systems, and machine learning communities, software defect datasets have seen significant growth in both supply and demand in recent years. This article enables users to identify datasets that best align with their research and practical needs, and it gives future dataset creators guidance on best practices for dataset development and maintenance. Based on the discussions of existing datasets, we also propose several future directions to improve the quality and accessibility of software defect datasets. To facilitate exploration and selection of software defect datasets, we provide an interactive website where users can browse and filter datasets based on their specific needs: https://defect-datasets.github.io/.

### v2

@@ 6. Uses of Software Defect Datasets (RQ4)
Understanding how software defect datasets are used in practice is essential not only for assessing their impact on research and practice, but also for revealing current needs. To this end, we conduct a systematic analysis of research work citing the top-five most-cited [^3] As of January 1, 2025, on Google Scholar. software defect dataset papers among the ones considered in this survey, which are Defects4J (DBLP:conf/issta/JustJE14), Bug Prediction Dataset (DBLP:journals/ese/DAmbrosLR12), Big-Vul (DBLP:conf/msr/FanL0N20), the ManyBugs and IntroClass benchmarks (DBLP:journals/tse/GouesHSBDFW15), and SWE-Bench (DBLP:conf/iclr/JimenezYWYPPN24). All five datasets are general-purpose and not limited to any specific application domain. Four datasets primarily focus on functional defects, while Big-Vul specifically targets security vulnerabilities. In terms of programming languages, Defects4J and the Bug Prediction Dataset collect defects from Java projects; Big-Vul and the ManyBugs and IntroClass benchmarks focus on C/C++ defects; and SWE-Bench, the most recent among them, is in Python.
@@
We categorize the uses of software defect datasets by conducting a systematic analysis of papers citing these datasets. Initially, we manually review a subset of citing papers to develop a preliminary categorization. Using an open-coding approach, we continuously refine the categories as we review additional papers. This taxonomy construction process involves over 150 randomly selected citing papers, leading to the following categorization scheme:
@@
The “empirical research” and “technical evaluations” categories are not mutually exclusive, as a paper may use a dataset for both purposes. For instance, a paper that empirically evaluates existing program repair tools using established software defect datasets is classified into both empirical research and the repair category of technical evaluations, while a paper that introduces a new repair technique and evaluates it falls solely under technical evaluation.
Based on the above scheme, we use the o3-mini model from OpenAI to automatically categorize papers using their titles and abstracts. The model is instructed with a structured prompt [^4] The full prompt is provided in the supplementary materials (Zhu et al., 2025). that enumerates our taxonomy and labeling rules, including that a paper may have multiple labels, that new benchmark is mutually exclusive with other categories, and that ambiguous cases should be marked as others. Two authors jointly validate a random sample of 100 LLM-labeled papers, independent of the open-coding process, since open coding did not record per-paper labels. All 100 sampled papers have satisfactory classifications: 93 exactly match the human labels, and only 7 show minor discrepancies (e.g., non-contradictory extra categories). For instance, paper (Xia and Zhang, 2024) is labeled as both “Empirical Study” and “Repair” by the model, whereas the human label includes only “Repair”, because the abstract highlights substantial empirical evaluation. Likewise, paper (Ding et al., 2025) is labeled as both “Empirical Study” and “Detection” for the same reason, while the human label includes only “Detection”. We have also inspected the model-provided reasoning summaries, which generally (i) propose candidate categories based on cues in the title and abstract, (ii) apply the taxonomy rules to include or exclude labels, and (iii) resolve overlaps by prioritizing the primary contribution over supporting components.
Figure 10(a) illustrates that empirical research and technical evaluations are the primary use categories for the citing papers analyzed across the five most-cited datasets. Technical evaluations, which can be further divided into evaluations of detection, testing, and repair techniques, demonstrate a growing demand for software defect datasets, as shown in Figure 10(b). This section focuses on these two primary categories to provide a comprehensive overview of the uses of the most-cited software defect datasets. As our analysis specifically focuses on direct dataset utilization, we exclude papers that introduce new benchmarks and the “others” category from the detailed discussion.
6.1. Empirical Research
Software defect datasets can serve as valuable resources for empirical research to understand various aspects of software defects. Such empirical research can be broadly categorized into two groups. On one hand, studies may focus directly on analyzing software defects themselves, typically with regard to their manifestations, symptoms, root causes, and fixes. For example, Sobreira et al. (2018) present an in-depth analysis of 395 patches from Defects4J. By analyzing the patch size, location, and the underlying repair patterns, the study provides insights and future direction for automated program repair, such as highlighting the necessity for repair strategies capable of effectively handling method calls. On the other hand, empirical research can also explore interactions between software defects and existing software engineering techniques. For example, Durieux et al. (2019) conduct an empirical evaluation on eleven automated program repair tools using five software defect datasets. Their study shows that the effectiveness of automated program repair tools varies significantly across different datasets and concludes various causes of failures for the automated program repair. Both types of empirical research significantly benefit from pre-collected, well-structured software defect datasets, as these datasets offer diverse, realistic, and accessible instances of software defects, thus enabling rigorous analysis and reliable comparisons across different studies.
@@ 6.2. Technical Evaluations
Another primary use of the most-cited software defect datasets is to evaluate existing or new software engineering techniques, such as fault detection/localization (e.g., (Sarhan et al., 2022; An and Yoo, 2022; Schlichtig et al., 2023; Zeng et al., 2022; Yan et al., 2022; Wen et al., 2021; Wang et al., 2021; Yan et al., 2021; Sarhan et al., 2021; Perez et al., 2021; Kim et al., 2021; Nilizadeh et al., 2024; Sun et al., 2024)), test generation/prioritization (e.g., (Zhang et al., 2024d; Jabbar et al., 2023; Rotaru and Vescan, 2023; Chaim et al., 2023; Chen et al., 2023; Almulla and Gay, 2022; Biswas et al., 2022; Gazzola et al., 2022)), and automated program repair (e.g., (Ye et al., 2023; Liu et al., 2023; Tian et al., 2023a; Fan et al., 2023; Chakraborty et al., 2022)). Fault detection/localization aims to identify defects in software in an effective, efficient, and automated manner. Test generation/prioritization is about generating or prioritizing test cases to maximize the likelihood of detecting defects, while minimizing the cost of testing. Both detection and testing techniques have long attracted significant interest from the software engineering, programming languages, and other communities, as such techniques are essential for improving software quality and reliability. The goal of automated program repair (Le Goues et al., 2019) is to fix defective software without human intervention, which is a challenging task that has gained increasing attention in recent years. The evaluation of these techniques is crucial for understanding their effectiveness, efficiency, and limitations, and software defect datasets play a critical role in this process. By providing standardized, realistic, and reproducible instances of software defects, software defect datasets enable researchers and practitioners to evaluate their techniques in a controlled environment.
@@
Figure 10(a) illustrates that detection and repair techniques are the most frequently evaluated. Defects4J is used for evaluating all three types of techniques, and it is also the predominant dataset used for evaluating testing techniques. The Bug Prediction Dataset (DBLP:journals/ese/DAmbrosLR12) and Big-Vul (DBLP:conf/msr/FanL0N20) are primarily used for evaluating detection techniques, while both the ManyBugs and IntroClass benchmarks (DBLP:journals/tse/GouesHSBDFW15) and SWE-Bench (DBLP:conf/iclr/JimenezYWYPPN24) are used more extensively for evaluating repair techniques. Figure 10(b) illustrates the unprecedented growth in the number of proposed software engineering techniques, particularly in the area of automated program repair. This surge in recent years underscores the increasing importance and demand for high-quality software defect datasets.
RQ4 Summary Software defect datasets are primarily utilized in empirical studies and technical evaluations. Empirical work investigates defect characteristics, root causes, and fixes, while technical evaluations emphasize testing, detection, and automated repair. The most-cited datasets are frequently reused in software engineering research. Their adoption indicates a shift toward evaluating automated techniques like automated repair, test generation, and defect detection, highlighting the increasing demand for executable and large-scale datasets suitable for data-driven and LLM-based methods.
@@ 7.1. Creation of Software Defect Datasets
Though various methodologies for constructing defect datasets have been proposed (Section 4), the balance between scale, diversity, quality, and cost remains a challenge. Intuitively, datasets of high quality tend to be smaller and more costly to create, such as the hand-crafted Defects4J (DBLP:conf/issta/JustJE14). In this sense, further efforts are still needed for fully automated defect mining. However, an important bottleneck for automated mining is the lack of a ground truth that indicates the actual defects present in the codebase. For example, keyword-based mining, the most commonly used technique, shows a relatively low precision of 51.9% and recall of 58.8% (Tian et al., 2012). It would be desirable to establish a ground truth without manual labelling or testing. BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19) and Bears (DBLP:conf/wcre/DelfimUMM19) have shown that by leveraging the fail-pass pattern in CI/CD pipelines, one can automatically mine defects with high precision. Such techniques may be adopted into the mining processes for other sources, without narrowing to CI/CD build histories. Similarly, ReposVul (DBLP:conf/icse/WangHGWC024) have shown that the combination of LLMs and static analysis can achieve up to 95% precision in defect mining. Such approaches can be further explored to improve the quality of defect datasets while keeping the manual effort low.
@@
Though tremendous efforts have been made in defect mining, the granularity of defect datasets is usually at the repository-, file-, or method-level, without fully precise defect location information that pinpoints defects, e.g., at the line-level. Ideally, datasets should isolate defects, e.g., by removing any unrelated code changes from a bug-fixing commit. For example, to isolate actual defects from newly added features, refactorings, and other code changes, Defects4J (DBLP:conf/issta/JustJE14) has manually reviewed and filtered the code diffs that fix a given bug. Such a manual review and filtering process is costly and time-consuming, though; thus, it is preferable to have automated techniques to precisely locate defects. Such isolation techniques could be based on refactoring detection (Jiang et al., 2021) or probabilistic delta debugging (DBLP:conf/kbse/SongWCCLLWP23). Another related problem is the annotation of defects, e.g., with the defect’s category. Though some datasets provide such annotations, they are typically manually labelled or inferred from other metadata such as exceptions (DBLP:conf/icse/DmeiriTWBLDVR19). Recently, machine learning and natural language processing techniques have shown promising results for automated defect annotation (DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/dsa/XuYWA19; Ahmed et al., 2024). More investigation is needed on the effectiveness and efficiency of such techniques, and on how they can be integrated into the process of creating defect datasets.
@@
Effectively organizing a dataset is crucial for maximizing its availability and usability. Currently, there is no standardized format for organizing software defect datasets, which can lead to difficulties in dataset access and usage. GitHub is the most popular platform for hosting datasets and tool chains, but it does not ensure long-term dataset availability and is susceptible to data loss or deletion. It is desirable to have datasets archived in a more stable and reliable repository, such as Zenodo or figshare, in addition to GitHub (Association for Computing Machinery (ACM), 2025). Furthermore, the organization of datasets should facilitate future evaluation of new techniques. Section 5.2 shows that half of the datasets provide only textual metadata (i.e., meta-level), which is insufficient for running software defects. Our analysis indicates that, after normalizing citation counts by dataset age (i.e., citations per year), execution-level datasets exhibit higher uptake (mean 53.65, median 12.00) than code-level (mean 21.01, median 12.75) and meta-level datasets (mean 18.07, median 8.50). Future efforts should focus on providing more comprehensive and reproducible software artifacts for defects. At a minimum, we recommend the following artifact reproducibility checklist: (1) a metadata index that records each defect and links it to the corresponding buggy and fixed code versions; (2) an environment descriptor, such as a container image or environment specification; (3) clear build and test instructions with expected outcomes; and (4) a runnable reproduction driver that demonstrates the defect and confirms the fix.
@@ 7.2. Maintenance of Software Defect Datasets
Ensuring the reproducibility of defects in software defect datasets is crucial for researchers to validate their findings and advance existing work. However, achieving reproducibility poses a significant challenge in software defect datasets, regardless of the construction methodology or dataset organization (Zhu and Rubio-González, 2023). Recent work shows that the major reason for defects not being reproducible is inaccessible dependencies, which datasets can mitigate by caching dependencies (Zhu and Rubio-González, 2023; DBLP:conf/icse/SaavedraSM24; DBLP:conf/msr/SilvaSM24). In this direction, more techniques are required to improve the long-term reproducibility of datasets while maintaining the balance between dataset size and usability.
@@
Software defect datasets can benefit from further evolution after initial publication, but most are not updated after release, leading to outdated information. Datasets that do not evolve may cause data leakage when evaluating machine learning-based approaches, especially LLMs, as defects may be in the training corpus (Tian et al., 2023b). Future datasets should support continuous updates by adding new defects while maintaining historical data for longitudinal studies. Continuous mining with automated pipelines (DBLP:conf/icse/DmeiriTWBLDVR19; Zhu et al., 2023) and community-driven efforts (McGuire et al., 2023) show promise for dataset evolution. However, frequent updates complicate comparisons with prior work, as different studies may use distinct dataset versions. Clear versioning, documentation, and stable baselines are crucial for fair and reproducible comparisons. Future research should explore best practices for balancing new defects with historical data maintenance and incentivizing dataset creators to evolve published datasets.
8. Threats to Validity
Selection and implementation bias
Despite our comprehensive search across four major bibliographic databases, we may miss work published outside these databases, or in venues not covered by our selection criteria. Also, excluding preprints and non-peer-reviewed works ensures quality but may miss emerging datasets. Automated scripts in data collection may contain unnoticed errors, and human judgment in screening may introduce subjectivity.
Annotation and interpretation bias
Manual annotation of dataset attributes involves subjective interpretation. Although we measure inter-rater agreement across multiple independent annotations, minor biases may still persist. The LLM-based analysis in Section 6 is manually validated, but some errors may remain due to model limitations.
Construct validity
Our analytical dimensions may omit factors such as community engagement, and differing classification granularity within each dimension may affect cross-study comparability. Our usage analysis prioritizes the most-cited datasets and may not generalize; dataset evolution after publication may also affect the stability of findings.
@@ 9. Conclusion
This survey provides a comprehensive overview of the landscape of software defect datasets, including their scope, construction, availability, usability, and practical uses. As fundamental resources to the software engineering, programming languages, computer security, computer systems, and machine learning communities, software defect datasets have seen significant growth in both supply and demand in recent years. This article enables users to identify datasets that best align with their research and practical needs, and it gives future dataset creators guidance on best practices for dataset development and maintenance. Based on the discussions of existing datasets, we also propose several future directions to improve the quality and accessibility of software defect datasets. To facilitate exploration and selection of software defect datasets, we make our artifact publicly available at (Zhu et al., 2025) and provide an interactive website where users can browse and filter datasets based on their specific needs: https://defect-datasets.github.io/.